OpenStudy (anonymous):

Jasmine wants to use her savings of $1128 to buy video games and music CDs. The total price of the music CDs she bought was $72. The video games cost $43 each. What is the maximum number of video games that Jasmine can buy with her savings?

OpenStudy (dogzcatz):

25

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with The table below shows two equations: Equation 1 |2x − 3| + 5 = 4 Equation 2 |5x + 3| − 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 2, −3.2. The solutions to equation 1 are x = 1, 2 and equation 2 has no solution. The solutions to equation 1 are x = 1, 2 and equation 2 has solutions x = 2, −3.2.